# Thermally activated particle motion in biased correlated Gaussian disorder potentials

### v1

@@ II Optimal Fluctuation Method
$\kappa\left(x-x^{\prime}\right)=\left\langle V\left(x\right)V\left(x^{\prime}\right)\right\rangle\,.$ (7)
@@
In the presence of a uniform bias field, $E=\text{const}$, the potential $V(x)$ can be represented as
@@ II.2.1 Monotonically decreasing covariance
As a simple but instructive example, let us suppose that the covariance decays as a power law, $\kappa(z\to\infty)\simeq B\,z^{-\alpha}$, where $\alpha>0$. Then Eq. (31) yields
@@ III.1 Simulation method
Instead of the poorly scalable matrix multiplications, a more efficient way to sample HGFs is to use the Circulant Embedding method Chan1994 ; Dietrich1997 . The method is based on embedding a covariance matrix $C_{i,j}$ of size $M\times M$ in a larger circulant covariance matrix $C^{\prime}_{ij}$ of size $M^{\prime}>2M$. Then, using the fact that a matrix multiplication with a circulant matrix implements convolution, one can replace the matrix equation (48) with multiplication in Fourier space. Therefore the required data (like the result of matrix diagonalization) can be easily pre-computed numerically by using the fast Fourier transform (FFT), while sampling a HGF requires only $\mathcal{O}(M^{\prime}\ln M^{\prime})$ operations.
@@
Here we are interested in extremely small probability densities which are virtually impossible to reach with conventional Monte Carlo (MC) simulations employing the Metropolis-Hastings algorithm. Therefore, in order to reach the large-$\Delta V$ tail of the distribution $\mathcal{P}(\Delta V)$ we employed the Wang-Landau (WL) algorithm WL1 ; WL2 . In a nutshell, the WL algorithm aims at estimating the density of states $\mathcal{P}(\Delta V)=\exp[-S(\Delta V)]$ on a chosen interval $a\leq\Delta V\leq b$, updating on each MC step the histogram of visited states $H(\Delta V)$ and adjusting the guess function $S(\Delta V)$ in an iterative manner, see Ref. Landau_Binder for details. That is, unlike the ordinary Metropolis-Hastings algorithm, where the acceptance/rejection decisions are Markovian, the Wang-Landau algorithm takes into account previously visited states.
@@ Acknowledgments
We are grateful to Satya N. Majumdar for a useful discussion. This research was supported by the Israel Science Foundation (Grant No. 1499/20).

### v2

@@ II Optimal Fluctuation Method
$\kappa\left(x-x^{\prime}\right)=\left\langle V\left(x\right)V\left(x^{\prime}\right)\right\rangle-\langle V(x)\rangle\langle V(x^{\prime})\rangle\,.$ (7)
@@
In the presence of a uniform bias field, $E=\text{const}$, we have $\langle V(x)\rangle=-Ex$, and the potential $V(x)$ can be represented as
@@ II.2.1 Monotonically decreasing covariance
As a simple and useful example, let us suppose that the covariance decays as a power law, $\kappa(z\to\infty)\simeq B\,z^{-\alpha}$, where $\alpha>0$. Then Eq. (31) yields
@@ III.1 Simulation method
Instead of the poorly scalable matrix multiplications, a more efficient way to sample HGFs is to use the Circulant Embedding method (CEM) Chan1994 ; Dietrich1997 . The method is based on embedding a covariance matrix $C_{i,j}$ of size $M\times M$ in a larger circulant covariance matrix $C^{\prime}_{ij}$ of size $M^{\prime}>2M$. Then, using the fact that a matrix multiplication with a circulant matrix implements convolution, one can replace the matrix equation (48) with multiplication in Fourier space. Therefore the required data (like the result of matrix diagonalization) can be easily pre-computed numerically by using the fast Fourier transform (FFT), while sampling a HGF requires only $\mathcal{O}(M^{\prime}\ln M^{\prime})$ operations. Some instructive implementations of the CEM for large deviation simulations of fractal Brownian motion (whose time derivative is a stationary Gaussian process) can be found in Refs. Hartmann_2 ; Hartmann_1 .
@@
Here we are interested in extremely small probability densities, which are virtually impossible to reach with conventional Monte Carlo (MC) simulations employing the Metropolis-Hastings algorithm. Therefore, in order to reach the large-$\Delta V$ tail of the distribution $\mathcal{P}(\Delta V)$, we employed the Wang-Landau (WL) algorithm WL1 ; WL2 . Unlike the ordinary Metropolis-Hastings algorithm, where the acceptance/rejection decisions are Markovian, the Wang-Landau algorithm takes into account information about previously visited states in such a way that it “forces” the algorithm to explore the available configurational space more quickly.
In a nutshell, the WL algorithm aims at estimating the density of states $\mathcal{P}(\Delta V)=\exp[-S(\Delta V)]$ on a chosen interval $a\leq\Delta V\leq b$, updating on each MC step the histogram of visited states $H(\Delta V)$ and adjusting the entropy function $S(\Delta V)$ in an iterative manner, see Ref. Landau_Binder for details.
At the start of the simulation, the histogram is initialized to zero, $H(\Delta V)=0$, and the entropy is set to some guess function (we use $S(\Delta V)=1$). Let $\bigl{\{}\vec{\xi}_{\text{r}},\vec{V}_{\text{r}},\Delta V_{\text{r}}\bigl{\}}$ represent the running configurations of the random vector, the disorder potential computed using the CEM, and the maximal potential barrier of $\vec{V}_{\text{r}}$, respectively. A proposed configuration of the random vector $\vec{\xi}_{\text{p}}$ is generated by changing a randomly chosen component $[\vec{\xi}_{\text{r}}]_{j}$ of the random vector according to the Gaussian distribution centered at $[\vec{\xi}_{\text{r}}]_{j}$:
$[\vec{\xi}_{\text{p}}]_{j}\sim\exp[-(x-[\vec{\xi}_{\text{r}}]_{j})^{2}/2].$ (49)
Then, the proposed configuration of the disorder potential $\vec{V}_{\text{r}}$ and its maximal potential barrier $\Delta V_{\text{r}}$ are computed using the CEM and the proposed random vector $\vec{\xi}_{\text{p}}$.
Every decision on whether to accept (a) or reject (r) the proposed configuration $\bigl{\{}\vec{\xi}_{\text{p}},\vec{V}_{\text{p}},\Delta V_{\text{p}}\bigl{\}}$ is made according to the transition probability
$p_{\text{a/r}}=\min\left(r(\vec{V}_{\text{p}}|\vec{V}_{\text{r}})\frac{\exp[-S(\Delta V_{\text{r}})]}{\exp[-S(\Delta V_{\text{p}})]},1\right),$ (50)
where $\text{where }r(\vec{V}_{\text{p}}|\vec{V}_{\text{r}})=\exp\left[\frac{\left[\vec{\xi}_{\text{r}}]_{j}^{2}-[\vec{\xi}_{\text{p}}\right]_{j}^{2}}{2}\right]\,.$
If the proposed configuration is rejected, the running configuration $\bigl{\{}\vec{\xi}_{\text{r}},\vec{V}_{\text{r}},\Delta V_{\text{r}}\bigl{\}}$ is kept. Following each decision, the action and the histogram are updated
$S(\Delta V_{\text{r}})\to S(\Delta V_{\text{r}})-f\,,$ (51)
$H(\Delta V_{\text{r}})\to H(\Delta V_{\text{r}})+1\,,$
where $f$ is a modification factor (initially, we set $f=1$). This process is repeated until the histogram of visited states $H(\Delta V)$ is sufficiently flat. As a measure of flatness, we used the condition
$0.9\times\overline{H(\Delta V)}\leq\min(H(\Delta V)),$ (52)
where $\overline{H(\Delta V)}$ is the mean value of the histogram. Once this condition is met, $f$ is reduced, $f\to f/2$, and the histogram is reset to zero. Then, the process continues until $f$ is sufficiently small.
It is known that, at early stages of a simulation, the Wang-Landau algorithm violates the detailed balance condition. Asymptotically, however, the detailed balance is recovered as the modification factor $f$ tends to zero Landau_Binder . We stopped the Wang-Landau simulations after 18 reductions of the modification factor $f$, which suffices for our purposes of large deviation simulations. This is because the statistical errors of the Wang-Landau algorithm, known to saturate Zhou , are of the order $\mathcal{O}(1)$, which is much smaller than the resulting entropy in our simulations, $S(\Delta V)\sim\mathcal{O}(10^{3})$.
@@ Acknowledgments
We are grateful to A. K. Hartmann and S. N. Majumdar for useful discussions. This research was supported by the Israel Science Foundation (Grant No. 1499/20).